# DTG : Diffusion-based Trajectory Generation for Mapless Global Navigation

### v2

@@ Abstract
We present a novel end-to-end diffusion-based trajectory generation method, DTG, for mapless global navigation in challenging outdoor scenarios with occlusions and unstructured off-road features like grass, buildings, bushes, etc. Given a distant goal, our approach computes a trajectory that satisfies the following goals: (1) minimize the travel distance to the goal; (2) maximize the traversability by choosing paths that do not lie in undesirable areas. Specifically, we present a novel Conditional RNN(CRNN) for diffusion models to efficiently generate trajectories. Furthermore, we propose an adaptive training method that ensures that the diffusion model generates more traversable trajectories. We evaluate our methods in various outdoor scenes and compare the performance with other global navigation algorithms on a Husky robot. In practice, we observe at least a $15\%$ improvement in traveling distance and around a $7\%$ improvement in traversability.
@@ III-A Problem Definition
Perceptual sensors used in our approach include a 3D LiDAR and the robot’s odometer. We utilize $C_{l}$ consecutive frames of LiDAR perception, $\mathcal{O}_{l}$, to capture the static and dynamic information of the environment in the robot’s vicinity. The odometer provides the latest $C_{v}$ consecutive frames of the robot’s velocities, $\mathcal{O}_{v}$, to encode the robot’s dynamic status. In this global navigation task, the goal is denoted as $g\in\mathcal{O}_{g}$. Thus, the observation contains $\mathcal{O}=\left\{\mathcal{O}_{l},\mathcal{O}_{v},\mathcal{O}_{g}\right\}$. Given the observation $\mathcal{O}$, our trajectory generator $DTG_{\theta}$ generates a trajectory $\bm{\tau}=\{\mathbf{w}_{1},...,\mathbf{w}_{M}\}$ that both satisfies traversability constraints and has minimum travel distance to the goal $g$. Here, $\mathbf{w}_{m}=\{x_{m},y_{m}\}$ represents the waypoints in the generated trajectory, which contains M waypoints in total starting from the robot. We define the travel distance of the trajectory $\bm{\tau}$ as the length of the shortest path from the last waypoint $\mathbf{w}_{M}$ of $\bm{\tau}$ to the target $g$. For each observation $o\in\mathcal{O}$, we have $DTG_{\theta}(\mathbf{o})=\bm{\tau}$. The problem can be formulated as the following equation:
@@ III-B2 Diffusion Model
The Diffusion model generates high-quality data by progressively denoising a Gaussian noise to some target data [ho2020denoising, song2020denoising], and it shows promising capabilities in robotics tasks [chi2023diffusion]. In our approach, the diffusion model takes the conditional vector $\mathbf{c}$ from the Perception Encoder for each observation and denoise a Gaussian distribution to a trajectory $\hat{\bm{\tau}}$. In recent years, multiple diffusion training strategies have been proposed, including DDPM [ho2020denoising], DDIM [song2020denoising], and VDM [vdm]. Because VDM [vdm] shows higher quality in data generation and better convergence than others, in this approach, we use the VDM [vdm] strategy to generate trajectories. In the diffusion process, neural networks (diffusion cells) are trained to learn noise or noisy data, and these two procedures are mathematically equivalent. Because of faster convergence and smaller converged loss, we use the diffusion cells to predict trajectories in our approach, DTG. Our diffusion generator contains N steps. As shown in Figure 2, from the final output ($\mathbf{x}_{0}$) to the Gaussian noise ($\mathbf{x}_{N}$), noise is introduced in each step, resulting in a noisy trajectory denoted as $\tilde{\mathbf{x}}_{t}=\sqrt{\alpha_{t}}\mathbf{x}_{t-1}+\sqrt{1-\alpha_{t}}\epsilon$, where $\alpha_{t}$ is the noise ratio at step $t$, which ranges from 1 to N. $\epsilon$ represents the noise itself. The predicted trajectory in step $t$ is $\hat{\mathbf{x}}_{t-1}=R_{\theta}(\tilde{\mathbf{x}}_{t},t)$, where $R_{\theta}(\cdot)$ is a diffusion cell. Because the U-Net [ho2020denoising, chi2023diffusion] is very computationally enhancive, we propose a novel diffusion cell to reduce the computational cost in Figure 2. Since we need the generated trajectories to remain always within traversable areas, we require the diffusion cell to also integrate environmental information. Inspired by [chi2023diffusion], we propose a Conditional RNN (CRNN), shown on the right side of Figure 2, that takes environmental information vector $\mathbf{c}$ and the step number $t$ as conditions for the generator, denoted by $R_{\theta}(\tilde{\mathbf{x}}_{t-1},t,\mathbf{c})$. Thus, we have the CRNN cell:
@@ III-C Training Strategy
$\mathcal{L}_{d}=\underset{t,\epsilon_{d}}{\mathbb{E}}\|\hat{\bm{\tau}}_{t}-s(\bm{\tau}_{gt},t)\|^{2}_{2}.$ (9)
@@ III-C2 Adaptive Training of Diffusion Models
where $d(\cdot)$ calculates the distance between the current waypoint $w_{m}$ in $\hat{\bm{\tau}}_{t}$ and its nearest non-traversable area The distance function clips the value to $[0-1]$ meters. We sample t from the diffusion step buffer $\mathbf{b}_{t}$, which stores the currently available diffusion steps. As shown in Algorithm 1, $i$ is the index of training epochs and $N_{t}=10$ is the maximum step number to apply to the traversability loss. $l_{d}(t-1)$ calculate the average of $5$ last recorded diffusion loss values at step $t-1$. If the average loss value is smaller than the threshold $H_{d}$, we add this step to the step buffer $\mathbf{b}_{t}$. We increase the range of diffusion steps incrementally to adapt the training of traversability to the training of ground truth trajectories. Finally, we have the total loss function for DTG: $\mathcal{L}=\mathcal{L}_{d}+\beta\mathcal{L}_{t}$, where $\beta$ is a hyperparameter.
@@ IV-A Implementation
As described in Section III, our major perceptual sensor is a 3Hz 3D Velodyne Lidar (VLP-16) with 16 channels. The inputs of this approach contain $C_{l}=3$ consecutive frames of Lidar and $C_{v}=20$ consecutive frames of historic velocities. The goal is set by converting the GPS value to meters. During training, the goals are randomly selected within 60 meters; for testing, the goals are selected beyond 50 meters. The trajectory generator, DTG, generates $M=16$ waypoints in each trajectory. The voxelization radius of the PointCNN in the Perception Encoder is 0.08m in this approach. The training data is the same as MTG [liang2023mtg], which is collected by a Husky robot. Velocities are collected from a 10Hz odometer. To keep all the perceptive information in the same time period, we select consecutive 10 frames of velocities as input. During training, the ground truth trajectories, generated by the A* algorithm with the shortest travel distance to the goal, are thresholded 15 meters from the robot’s position, and have 16 waypoints each. The training and evaluation are processed in a computer with an NVIDIA RTX A5000 GPU and an Intel Xeon(R) W-2255 CPU, and the real-world experiment is executed on a laptop with an Intel i7 CPU and one Nvidia GTX 1080 GPU. In the real-world experiment, the diffusion model generates trajectories in 5Hz. The network details of the architecture are in Appendix VI-C[supplements].
@@ IV-B Evaluation
Traversability: Given a trajectory $\hat{\bm{\tau}}$, the traversability is calculated as Equation 11. The $c(\cdot)$ tells if the waypoint $\mathbf{w}_{m}$ is in the traversable area $\mathcal{A}$. For K secnearios, we have the average traversability of the approach: $\frac{1}{K}\sum_{i=1}^{K}tr(\mathcal{A},\hat{\bm{\tau}})$.
@@
ViNT 66.02 79.02 0.69 113.49
NoMaD 64.54 77.86 0.24 72.67
MTG 80.78 83.11 0.01 101.61
CVAE 92.26 85.89 0.01 113.96
DTG/t 93.12 85.57 0.13 128.38
@@
Comparisons: This section shows the efficacy of DTG for mapless outdoor global navigation. From Table I, we observe our approach, DTG, outperforms other SOTA approaches. Specifically, we achieve by at least $15\%$ in distance ratio and $12.6\%$ in traversability improvement compared with ViNT and NoMaD. NoMaD and ViNT only take RGB images, so the robot is not very robust in generating trajectories only in traversable areas. NoMaD is slightly better than ViNT w.r.t. traversability and inference time. Our approach outperforms MTG around $7\%$ in traversability and $16\%$ in distance ratio. The MTG has relatively good traversability but cannot choose the trajectory with the best distance ratio because in MTG the trajectories are generated by only comparing the straight distance to the goal instead of estimating the real travel distance of the trajectory. DTG also outperforms NoMaD and ViNT in inference time by 0.11s and 0.56s, but the model size is bigger than ViNT, NoMaD, and MTG.

### v3

@@ Abstract
We present a novel end-to-end diffusion-based trajectory generation method, DTG, for mapless global navigation in challenging outdoor scenarios with occlusions and unstructured off-road features like grass, buildings, bushes, etc. Given a distant goal, our approach computes a trajectory that satisfies the following goals: (1) minimize the travel distance to the goal; (2) maximize the traversability by choosing paths that do not lie in undesirable areas. Specifically, we present a novel Conditional RNN(CRNN) for diffusion models to efficiently generate trajectories. Furthermore, we propose an adaptive training method that ensures that the diffusion model generates more traversable trajectories. We evaluate our methods in various outdoor scenes and compare the performance with other global navigation algorithms on a Husky robot. In practice, we observe at least a $15\%$ improvement in traveling distance and around a $7\%$ improvement in traversability. Video and Code: https://github.com/jingGM/DTG.git.
@@ III-A Problem Definition
Perceptual sensors used in our approach include a 3D LiDAR and the robot’s odometer. We utilize $C_{l}$ consecutive frames of LiDAR perception, $\mathcal{O}_{l}$, to capture the static and dynamic information of the environment in the robot’s vicinity. The odometer provides the latest $C_{v}$ consecutive frames of the robot’s velocities, $\mathcal{O}_{v}$, to encode the robot’s dynamic status. In this global navigation task, the goal is denoted as $g=\left\{g_{x},g_{y}\right\}\in\mathcal{O}_{g}$. Thus, the observation contains $\mathcal{O}=\left\{\mathcal{O}_{l},\mathcal{O}_{v},\mathcal{O}_{g}\right\}$. Given the observation $\mathcal{O}$, our trajectory generator $DTG_{\theta}$ generates a trajectory $\bm{\tau}=\{\mathbf{w}_{1},...,\mathbf{w}_{M}\}$ that both satisfies traversability constraints and has minimum travel distance to the goal $g$. Here, $\mathbf{w}_{m}=\{x_{m},y_{m}\}$ represents the waypoints in the generated trajectory, which contains M waypoints in total starting from the robot. We define the travel distance of the trajectory $\bm{\tau}$ as the length of the shortest path from the last waypoint $\mathbf{w}_{M}$ of $\bm{\tau}$ to the target $g$. For each observation $o\in\mathcal{O}$, we have $DTG_{\theta}(\mathbf{o})=\bm{\tau}$. The problem can be formulated as the following equation:
@@ III-B2 Diffusion Model
The Diffusion model generates high-quality data by progressively denoising a Gaussian noise to some target data [ho2020denoising, song2020denoising], and it shows promising capabilities in robotics tasks [chi2023diffusion]. In our approach, the diffusion model takes the conditional vector $\mathbf{c}$ from the Perception Encoder for each observation and denoise a Gaussian distribution to a trajectory $\hat{\bm{\tau}}$. In recent years, multiple diffusion training strategies have been proposed, including DDPM [ho2020denoising], DDIM [song2020denoising], and VDM [vdm]. Because VDM [vdm] shows higher quality in data generation and better convergence than others, in this approach, we use the VDM [vdm] strategy to generate trajectories. In the diffusion process, neural networks (diffusion cells) are trained to learn noise or noisy data, and these two procedures are mathematically equivalent. Because of faster convergence and smaller converged loss, we use the diffusion cells to predict trajectories in our approach, DTG. Our diffusion generator contains N steps. As shown in Figure 2, from the final output ($\mathbf{x}_{0}$) to the Gaussian noise ($\mathbf{x}_{N}$), noise is introduced in each step, resulting in a noisy trajectory denoted as $\tilde{\mathbf{x}}_{t}=\sqrt{\alpha_{t}}\mathbf{x}_{t-1}+\sqrt{1-\alpha_{t}}\epsilon$, where $\alpha_{t}$ is the noise ratio at step $t$, which ranges from 1 to N. $\epsilon$ represents the noise itself. The predicted trajectory in step $t$ is $\hat{\mathbf{x}}_{t-1}=R_{\theta}(\tilde{\mathbf{x}}_{t},t)$, where $R_{\theta}(\cdot)$ is a diffusion cell. Because the U-Net [ho2020denoising, chi2023diffusion] is very computationally intensive, we propose a novel diffusion cell to reduce the computational cost in Figure 2. Since we need the generated trajectories to remain always within traversable areas, we require the diffusion cell to also integrate environmental information. Inspired by [chi2023diffusion], we propose a Conditional RNN (CRNN), shown on the right side of Figure 2, that takes environmental information vector $\mathbf{c}$ and the step number $t$ as conditions for the generator, denoted by $R_{\theta}(\tilde{\mathbf{x}}_{t-1},t,\mathbf{c})$. Thus, we have the CRNN cell:
@@ III-C Training Strategy
$\mathcal{L}_{d}=\underset{t,\epsilon_{d}}{\mathbb{E}}\|\hat{\bm{\tau}}_{t}-\bm{\tau}_{gt}\|^{2}_{2}.$ (9)
@@ III-C2 Adaptive Training of Diffusion Models
where $d(\cdot)$ calculates the distance between the current waypoint $w_{m}$ in $\hat{\bm{\tau}}_{t}$ and its nearest non-traversable area The distance function clips the value to $[0-1]$ meters. We sample t from the diffusion step buffer $\mathbf{b}_{t}$, which stores the currently available diffusion steps. As shown in Algorithm 1, $m_{\theta}(\cdot)$ is the neural network model and $\theta$ represents the parameters. $i$ is the index of training epochs and $N_{t}=10$ is the maximum step number to apply to the traversability loss. $l_{d}(t-1)$ calculate the average of $5$ last recorded diffusion loss values at step $t-1$. If the average loss value is smaller than the threshold $H_{d}$, we add this step to the step buffer $\mathbf{b}_{t}$. We increase the range of diffusion steps incrementally to adapt the training of traversability to the training of ground truth trajectories. Finally, we have the total loss function for DTG: $\mathcal{L}=\mathcal{L}_{d}+\beta\mathcal{L}_{t}$, where $\beta$ is a hyperparameter.
@@ IV-A Implementation
As described in Section III, our major perceptual sensor is a 3Hz 3D Velodyne Lidar (VLP-16) with 16 channels. The inputs of this approach contain $C_{l}=3$ consecutive frames of Lidar and $C_{v}=20$ consecutive frames of historic velocities. The goal is set by converting the GPS value to meters. During training, the goals are randomly selected within 60 meters; for testing, the goals are selected beyond 50 meters. The trajectory generator, DTG, generates $M=16$ waypoints in each trajectory. The voxelization radius of the PointCNN in the Perception Encoder is 0.08m in this approach. The training data is the same as MTG [liang2023mtg], which is collected by a Husky robot. The training and evaluation datasets are in different areas as shown in [liang2023mtg]. Velocities are collected from a 10Hz odometer. To keep all the perceptive information in the same time period, we select consecutive 10 frames of velocities as input. During training, the ground truth trajectories, generated by the A* algorithm with the shortest travel distance to the goal, are thresholded 15 meters from the robot’s position, and have 16 waypoints each. The training and evaluation are processed in a computer with an NVIDIA RTX A5000 GPU and an Intel Xeon(R) W-2255 CPU, and the real-world experiment is executed on a laptop with an Intel i7 CPU and one Nvidia GTX 1080 GPU. In the real-world experiment, the diffusion model generates trajectories in 5Hz. The network details of the architecture are in Appendix VI-C[supplements].
@@ IV-B Evaluation
Traversability: Given a trajectory $\hat{\bm{\tau}}$, the traversability is calculated as Equation 11. The $c(\cdot)$ tells if the waypoint $\mathbf{w}_{m}$ is in the traversable area $\mathcal{A}$. For K scenarios, we have the average traversability of the approach: $\frac{1}{K}\sum_{i=1}^{K}tr(\mathcal{A},\hat{\bm{\tau}})$.
@@
ViNT RGB 66.02 79.02 0.69 113.49
NoMaD RGB 64.54 77.86 0.24 72.67
MTG Lidar 80.78 83.11 0.01 101.61
CVAE Lidar 92.26 85.89 0.01 113.96
DTG/t Lidar 93.12 85.57 0.13 128.38
@@
Comparisons: This section shows the efficacy of DTG for mapless outdoor global navigation. From Table I, we observe our approach, DTG, outperforms other SOTA approaches. Specifically, we achieve by at least $15\%$ in distance ratio and $12.6\%$ in traversability improvement compared with ViNT and NoMaD. NoMaD and ViNT only take RGB images, so the robot is not very robust in generating trajectories only in traversable areas. NoMaD is better than ViNT w.r.t. inference time but slightly worse in traversability. Our approach outperforms MTG around $7\%$ in traversability and $16\%$ in distance ratio. The MTG has relatively good traversability but cannot choose the trajectory with the best distance ratio because in MTG the trajectories are generated by only comparing the straight distance to the goal instead of estimating the real travel distance of the trajectory. DTG also outperforms NoMaD and ViNT in inference time by 0.11s and 0.56s, but the model size is bigger than ViNT, NoMaD, and MTG.
@@ V Conclusion, Limitations, and Future Work
ACKNOWLEDGMENT
This work was supported in part by ARO Grants W911NF2310046, W911NF2310352 and U.S. Army Cooperative Agreement W911NF2120076